# Prediction Sets and Conformal Inference with

### v2

@@ Abstract
Given data on a scalar random variable $Y$, a prediction set for $Y$ with miscoverage level $\alpha$ is a set of values for $Y$ that contains a randomly drawn $Y$ with probability $1-\alpha$, where $\alpha\in(0,1)$. Among all prediction sets that satisfy this coverage property, the oracle prediction set is the one with the smallest volume. This paper provides methods for inference on such prediction sets given observed conditioning covariates when $Y$ is censored or measured in intervals. We first characterise the oracle prediction set under interval censoring and develop a consistent estimator for the shortest prediction interval that satisfies this coverage property. We then extend these consistency results to accommodate cases where the prediction set consists of multiple disjoint intervals. Second, we use conformal inference to construct a prediction set that achieves a particular notion of finite-sample validity under censoring and maintains consistency as sample size increases. This notion exploits exchangeability to obtain finite sample guarantees on coverage using a specially constructed conformity score function. The procedure accomodates the prediction uncertainty that is irreducible (due to the stochastic nature of outcomes), the modelling uncertainty due to partial identification and also sampling uncertainty that gets reduced as samples get larger. We conduct a set of Monte Carlo simulations and an application to data from the Current Population Survey. The results highlight the robustness and efficiency of the proposed methods.
@@ 1 Introduction
Interval data is pervasive. Surveys usually employ “bracketing” strategies to address item non-response. For example, respondents are often asked whether their money in savings accounts, in investments or their income fall in a sequence of brackets such as [\$$[\$10K,\$20K],[\$20K,\$50K]$ etc., while also given the option of providing a point answer to such questions (see heeringa1995unfolding and moore2001using). Interval censored data also accommodate often found patterns such as right censoring, where an outcome variable is greater than a specified value (i.e., $Y\in[Y^{L},\infty)$)); left censored data, where an outcome variable is less than a certain value (i.e., $Y\in(-\infty,Y^{U}])$); and data from competing risks, where one typically observes the $\max$ or $\min$ of variables of interest. Such data typically involve outcomes of interest such as income, wealth, unemployment duration and cause-specific mortality. This paper provides inference procedures for the prediction of a variable of interest $Y$ when we observe $[Y^{L},Y^{U}]\ni Y$ instead.
More specifically, suppose we have a random variable $Y$ defined on $\mathcal{Y}\subset\mathbb{R}$ with distribution $P$. One parameter of interest is a prediction set $C$ defined as the set of $y\in\mathcal{Y}$ such that $P(y\in C)\geq 1-\alpha$ for $\alpha\in(0,1)$ where we say that such a set $C$ has miscoverage level $\alpha\in(0,1).$ The oracle prediction set is the set with the smallest volume that has this (mis)coverage property and is related to the level set of the density of $Y$ (when this density exists). With covariates, we have access to a sample $(X_{i},Y_{i})\in\mathcal{X}\times\mathcal{Y}$, $i=1,\dots,n$, with $\mathcal{X}\subset\mathbb{R}^{d}$ and $\mathcal{Y}\subset\mathbb{R}$. We use these data to construct a prediction set $C_{n}(x)\equiv C(x;X_{1},Y_{1},\dots,X_{n},Y_{n})\subset\mathcal{Y}$ for a given miscoverage level $\alpha$ at $x\in\mathcal{X}$. It is desirable that suc prediction set $C_{n}$ is both consistent for the oracle prediction set and has a finite sample coverage property:
@@
where $\operatorname{P}$ is the distribution over $(X_{i},Y_{i})$ for $i=1,\dots,n+1$. This paper constructs such a set when $Y$ is interval censored. The exact coverage property that our construction provides is given in more details in Section 2 and Section 4 below.
@@
$N\left(X_{n+1}^{\prime}\hat{\beta},\ {\sigma}^{2}\left(1+X_{n+1}^{\prime}(\mathbf{X}^{\prime}\mathbf{X})^{-1}X_{n+1}\right)\right),$
@@
This “operational density” summarizes both the sampling uncertainty in estimating $\hat{\beta}$ and the disturbance uncertainty that comes from the intrinsic randomness in the conditional distribution $Y_{n+1}\mid X_{n+1}$. One can use it to derive a forecast density with a particular coverage property. In this paper, we use the recent theory on conformal inference to show formally that a particularly constructed predictive set ${C}_{n}$ has finite sample coverage property when $Y$ is censored and without parametric assumptions, while enjoying the asymptotic property of being consistent for the oracle prediction set under the partially identified true population distribution. This conformal inference procedure in turn requires the definition of particular score functions to handle interval censoring. This score is similar to ones used in the conformal literature but is here modified to handle censoring. In particular, the score function we use is negative when the constructed prediction set contains the interval outcomes, but is otherwise positive, and as a result accounts for both over- and under-coverage.
@@
In order to obtain the proposed conformal prediction set, we first characterise the oracle prediction sets under interval censoring and provide a feasible optimisation problem that relates the oracle prediction sets to the conditional distribution of the observed upper and lower brackts, $(Y^{L},Y^{U})$. We use a kernel based conditional distribution function estimator as a basis for the construction of a consistent set estimator for the oracle prediction set. Interestingly, the oracle prediction set may be the union of a collection of disjoint intervals in cases for instance where the prediction density is multimodal. We provide an approach that allows for the set estimator to be a union of disjoint intervals with the prescribed coverage. In cases with a multimodal density, a union of intervals can have a smaller volume than one interval.
@@
We then adapt recent work on conformal inference to propose a conformal prediction set using a sample of exchangeable data. This proposed construction is shown to have guarantees for finite sample coverage and to achieve asymptotic efficiency in terms of the volume of the prediction set. As a competing alternative prediction set estimator, we show in the appendix that aconformal quantile regression method provides a valid prediction set for interval censored outcomes which can be easier to implement in practice, but is in general not efficient in terms of the volume of the prediction set. In addition, quantile regression based methods are not suitable when the prediction set is not a single interval.
@@
Interval censoring may arise from fixed bracketing and random censoring from “unfolding brackets” or survey schemes, see heeringa1995unfolding and moore2001using. See also the bracketing approach to outcomes using the Heatlh and Retirement Survey in manski2002InferenceRegressions. We thus illustrate our contruction using a set of Monte Carlo experiments, where we consider both fixed and random censoring. Our proposed prediction set is shown to have adequate finite sample coverage properties and to have smaller volume compared with the alternative prediction sets constructed with quantile regression methods. In addition, we construct prediction sets using the Current Population Survey (CPS) data. Among people who report their income, a significant fraction (more than 20%) of the observations are interval censored. Interval censored observations are sometimes discarded or handled by imputation methods that rely on assumptions such as missing at random. Our proposed method prediction set is robust in the sense that we do not require a specification of the censoring mechanism and we demonstrate using a hold-out subset of the CPS data that our prediction set has better finite sample coverage of the interval censored incomes compared with prediction sets constructed with hot-deck imputation in place. Naturally, without relying on imputation algorithms, our oracle prediction sets are wider than ones obtained with uncensored outcomes, but our constructed sets are more robust and hence more trustworthy.
@@
The paper combines insights from the partial identification literature in econometrics and conformal inference results from the recent statistics literature. The partial identification literature in econometrics (see manski2003PartialIdentification) takes the view that one should try and make inferences under minimal assumptions, even if these assumptions do not allow for point identification of the parameter of interest. The motivation for such a view is that inferences are less credible and conclusions are suspect when these rely on non-testable or non credible assumptions. In our context for example, if we are interested in $\beta=E[Y]$ and we observe $[Y^{L},Y^{U}]$ rather than $Y$, then a partial identification approach would just say that, given the data, the identified set for $\beta$ (or all we can learn about $\beta$ given the data) is the interval $\left[E[Y_{0}],E[Y_{1}]\right].$ Imputation based inference uses models to predict a $Y^{P}$ within $[Y^{L},Y^{U}]$ and takes $E[Y^{P}]$ as a proxy for $\beta=E[Y]$. The credibility of this exercise depends on the imputation model, i.e. the set of assumptions it relies on. In the context of interval data, manski2002InferenceRegressions study identification of linear regression coefficients when the data (on regressors or outcomes) is interval censored using a partial identification approach. See tamer2010PartialIdentification, molinari2020ChapterMicroeconometrics, and kline2023RecentDevelopments for reviews on the developments in the partial identification literature.
@@
In order to obtain a prediction interval that has finite-sample coverage guarantees, we use the method of conformal inference, which is a distribution-free method that builds upon the idea of exchangeability and permutation tests (vovk2005AlgorithmicLearning, vovk2009LinePredictive). The idea of conformal inference has become increasingly popular in the machine learning literature since it requires no assumptions on the underlying distribution other than the samples being exchangeable and it also allows the use of any predictor that treats the data symmetrically and hence preserves exchangeability (barber2023ConformalPrediction discusses the validity of conformal inference when the data is not i.i.d. and the predictor is not symmetric). Conformal inference has been applied to obtain valid distribution-free prediction sets lei2013DistributionfreePrediction and has been extended to non-parametric regression lei2014DistributionfreePrediction and high-dimensional settings lei2018DistributionfreePredictive. romano2019ConformalizedQuantile considers conformal quantile regression to construct prediction intervals instead of conditional mean regression. sadinle2019LeastAmbiguous proposes a way to predict the label of $Y$. chernozhukov2021DistributionalConformal considers construction of valid prediction intervals with distributional regression method. We add to this literature by allowing the outcome observations to be interval-censored or interval-valued.
@@
The classical way to construct estimator for the oracle prediction set relies on the level set estimation, (see, for example, wilks1941DeterminationSample, polonik1995MeasuringMass and samworth2010AsymptoticsOptimal). The uniform consistency of the nonparametric estimator of the conditional distribution is an extension of classical results on the uniform convergence rate of the kernel-based density and regression estimators such as einmahl2000EmpiricalProcess, gine2002RatesStrong and hansen2008UniformConvergence. This paper is also related to the literature on quantile regression (see koenker2017QuantileRegression for a review).
@@
This paper is organised as follows. In Section 2, we provide the setup and definitions of optimal sets and approaches to obtaining estimators of such sets with censoring. Section 3 provides a study of consistency of such sets using kernel estimators of the conditional distribution. Section 4 provides results on finite sample coverage using a conformalized procedure under interval censoring. Section 5 provides some Monte Carlo evidence and Section 6 applies our methods to US CPS Data. Section 7 concludes. The proofs for the theorems in the main text as well as some further results and comments are collected in the Appendix.
@@ 2 Setup and definitions
Suppose there is a random sample $\bar{\mathcal{D}}=\{\bar{Z}_{i}=(X_{i},Y_{i},Y_{i}^{L},Y_{i}^{U})\in\mathbb{R}^{d+3}:i\in\mathcal{I}\}$ with sample size $\left\lvert\mathcal{I}\right\rvert=n$ drawn independently from a joint distribution $P$ [^1] In order to perform conformal inference, it is sufficient that the sample satisfies the weaker condition of exchangeability. Here independence is assumed for simplicity and for establishing the asymptotics for the nonparametric estimators we will use. barber2023ConformalPrediction studies conformal inference beyond exchangeability and is applicable for time series data. Our results can be readily extended to the time series settings combined with their procedure. We only observe the subvector $Z_{i}=(X_{i},Y_{i}^{L},Y_{i}^{U})$, and the latent outcome $Y_{i}$ is not directly observed but assumed to satisfy the condition $Y_{i}^{L}\leq Y_{i}\leq Y_{i}^{U}$. For each $i\in\mathcal{I}$, the predictor $X_{i}$ is a $d$-dimensional random vector with support $\mathcal{X}\subset\mathbb{R}^{d}$. The latent outcome $Y_{i}$ and the observed lower and upper bounds $Y_{i}^{L}$ and $Y_{i}^{U}$ are one-dimensional. Let $\bar{Z}=(X,Y,Y^{L},Y^{U})$ denote a generic sample from the joint distribution $P$. We will use notation such as $P_{Y}$ to denote the marginal distribution of $Y$, and $P_{Y\mid X}$ to denote the distribution of $Y$ conditional on $X$. Note here that the joint distribution $P_{X,Y^{L},Y^{U}}$ of $(X,Y^{L},Y^{U})$ is identified given the observed sample $\mathcal{D}=\left\{Z_{i}=(X_{i},Y_{i}^{L},Y_{i}^{U})\in\mathbb{R}^{d+2}:i\in\mathcal{I}\right\}$, but the joint distribution $P_{X,Y}$ of predictors $X$ and latent outcome $Y$ is not point identified given the observed sample $\mathcal{D}$ without restrictions on the censoring mechanism.
A prediction set is defined as a set $C\subset\mathcal{X}\times\mathcal{Y}$, with the interpretation that, given a realisation of the predictor $X=x$, we predict that $Y$ is likely to fall inside the section $C(x)=\left\{y:(x,y)\in C\right\}$. In the absence of interval censoring, estimation of a prediction set that satisfies certain optimality criteria (i.e. the minimal prediction set with coverage guarantee) has been studied in the literature; see lei2014DistributionfreePrediction, *barber2021LimitsDistributionfree among others. As we discuss later in this section, the oracle prediction set without interval censoring is related to the upper level set of the density function of $Y$ given $X$ under point identification.
@@
There are two criteria for an oracle prediction set: validity and efficiency [see][]vovk2009LinePredictive,lei2013DistributionfreePrediction. Let $\mathcal{P}_{I}$ be the identified set of the joint distribution of $(X,Y)$ under censoring. That is, $\mathcal{P}_{I}$ is the set of joint distributions of $(X,Y)$ that are compatible with the joint distribution $P_{X,Y^{L},Y^{U}}$ and some censoring mechanism. A prediction set $C$ is valid if it satisfies the following coverage properties for a new sample $(X_{n+1},Y_{n+1})$ drawn from the same distribution $P$ at a given miscoverage level $\alpha$.
@@
When the model is point identified, $\mathcal{P}_{I}=\left\{P_{X,Y}\right\}$ is a singleton set, the conditions in Equations (2) and (3) reduce to the standard conditions of validity, see for example, lei2014DistributionfreePrediction. We will simply refer to these two conditions as marginal validity and conditional validity henceforth.
It is natural to require a prediction set to be valid, so that we have confidence it will contain the true value of the outcome with a certain probability. However, the requirement of validity by itself is vacuous. For example, the prediction set $C$ such that $C(x)=(-\infty,\infty)$, for all $x\in\mathcal{X}$ is technically valid at any miscoverage level but it is also uninformative. Therefore, our target parameter of interest is an oracle prediction set, denoted by $C^{*}_{\mathcal{P}_{I}}$ which is defined as the minimal-volume prediction set that satisfies the validity conditions in Definition 2.1. For simplicity, we will omit the subscript and write $C^{*}=C^{*}_{\mathcal{P}_{I}}$ whenever the dependence on the identified set ${\mathcal{P}_{I}}$ is clear. Since the minimal-volume prediction set is intuitively the most informative among all valid prediction sets, this property is also called efficiency. [^2] This notion of efficiency differs from the usual definition, where an estimator is considered efficient if it has the smallest variance within a class of estimators. However, the concepts are related, as an estimator with smaller variance generally leads to the construction of tighter and more informative confidence intervals.
For the sake of measurability and interpretability, we will focus on the prediction sets such that $C(x)$ takes the form of a union of a (potentially unknown) number of disjoint intervals for each $x\in\mathcal{X}$. Specifically, we assume that $C(x)\in\mathcal{C}$, $P_{X}$-almost surely, where
@@
where $\mu$ denotes the Lebesgue measure. Since conditional validity implies marginal validity, the oracle predition set will also be marginally valid under partial identification.
Before proposing a feasible estimation procedure for the oracle prediction set $C^{*}$, we explain why the existing estimation methods are not applicable for prediction sets under point identification in the case of censoring. This analysis highlights the challenges posed by partial identification and the difference between our approach and the classical level set method studied in the literature.

### v3

@@ Abstract
Given data on a scalar random variable $Y$, a prediction set for $Y$ with miscoverage level $\alpha$ is a set of values for $Y$ that contains a randomly drawn $Y$ with probability $1-\alpha$, where $\alpha\in(0,1)$. Among all prediction sets that satisfy this coverage property, the oracle prediction set is the one with the smallest volume. This paper provides estimation methods of such prediction sets given observed conditioning covariates when $Y$ is censored or measured in intervals. We first characterise the oracle prediction set under interval censoring and develop consistent estimators for the oracle prediction intervals and prediction sets consisting of multiple disjoint intervals. We use conformal inference to construct a prediction set that achieves finite-sample validity under censoring and maintains consistency as sample size increases, using a conformity score function designed for interval data. The procedure accommodates the prediction uncertainty that is irreducible (due to the stochastic nature of outcomes), the modelling uncertainty due to partial identification and also sampling uncertainty that gets reduced as samples get larger. We conduct a set of Monte Carlo simulations and an application to data from the Current Population Survey. The results highlight the robustness and efficiency of the proposed methods.
Keywords: Oracle predictions sets, partial identification, nonparametric estimation, uniform consistency of set estimator.
@@ 1 Introduction
Interval data is pervasive. Surveys usually employ “bracketing” strategies to address item non-response. For example, respondents are often asked whether their money in savings accounts, in investments or their income fall in a sequence of brackets such as [\$$[\$10K,\$20K],[\$20K,\$50K]$ etc., while also given the option to provide a point answer to such questions (see Heeringa and Suzman, (1995) and Moore and Loomis, (2001)). Interval valued data also accommodate often found patterns such as right censoring, where an outcome variable is greater than a specified value (i.e., $Y\in[Y^{L},\infty)$)); left censored data, where an outcome variable is less than a certain value (i.e., $Y\in(-\infty,Y^{U}])$); and data from competing risks, where one typically observes the $\max$ or $\min$ of variables of interest. Such data typically involve outcomes of interest such as income, wealth, unemployment duration and cause-specific mortality. This paper provides inference procedures for the prediction of a variable of interest $Y$ when $Y$ is not observed, but we instead observe a sample of $[Y^{L},Y^{U}]$ such that $Y\in[Y^{L},Y^{U}]$.
More specifically, suppose we have a random variable $Y$ defined on $\mathcal{Y}\subset\mathbb{R}$ with distribution $P$. One parameter of interest is a prediction set $C$ defined as the set of $y\in\mathcal{Y}$ such that $P(y\in C)\geq 1-\alpha$ for $\alpha\in(0,1)$ where we say that such a set $C$ has miscoverage level $\alpha\in(0,1).$ The oracle prediction set is the set with the smallest volume that has this (mis)coverage property and is related to the level set of the density of $Y$ (when this density exists). With covariates, we have access to a sample $(X_{i},Y_{i})\in\mathcal{X}\times\mathcal{Y}$, $i=1,\dots,n$, with supports $\mathcal{X}\subset\mathbb{R}^{d}$ and $\mathcal{Y}\subset\mathbb{R}$. We use these data to construct prediction sets $C_{n}(x)\equiv C(x;X_{1},Y_{1},\dots,X_{n},Y_{n})\subset\mathcal{Y}$ for a given miscoverage level $\alpha$ at $x\in\mathcal{X}$. It is desirable that such prediction set $C_{n}$ is both consistent for the oracle prediction set and has a finite sample coverage property:
@@
where $\operatorname{P}$ is the distribution over $(X_{i},Y_{i})$ for $i=1,\dots,n+1$. This paper constructs such a set when $Y$ is interval valued. The exact finite-sample coverage property that our construction provides is given in more details in Section 4 .
@@
$N\left(X_{n+1}^{\prime}\hat{\beta},\ {\sigma}^{2}\left(1+X_{n+1}^{\prime}(\mathbf{X}^{\prime}\mathbf{X})^{-1}X_{n+1}\right)\right).$
This “operational density” summarizes both the sampling uncertainty in estimating $\hat{\beta}$ and the disturbance uncertainty that comes from the intrinsic randomness in the conditional distribution $Y_{n+1}\mid X_{n+1}$. One can use it to derive a forecast density with a particular coverage property.
The prediction set is closely related to quantile methods: the minimal volume prediction set would be given by $C(X_{n+1})\equiv[\tau_{0}(X_{n+1}),\tau_{1}(X_{n+1})]$, where $\tau_{0}(X_{n+1})$ and $\tau_{1}(X_{n+1})$ are appropriately chosen quantiles for $Y$ given $X$ so that one can obtain a coverage rate $(1-\alpha)$ at a minimum volume. Taking the Gaussian specification above, for example, with $\alpha=5\%$ and assuming $\beta$ to be known, $\tau_{0}(X_{n+1})=X_{n+1}^{\prime}\beta-1.96\sigma$, $\tau_{1}(X_{n+1})=X_{n+1}^{\prime}\beta+1.96\sigma$ where $1.96\sigma$ is the $95\%$-quantile for $|\varepsilon_{n+1}|\equiv|Y_{n+1}-X_{n+1}^{\prime}\beta|$. When $\sigma$ is known as presumed above, but $\beta$ is not, estimates for these can be obtained as $\hat{\tau}_{0}(X_{n+1})=X_{n+1}^{\prime}\hat{\beta}-1.96\sigma$ and $\hat{\tau}_{1}(X_{n+1})=X_{n+1}^{\prime}\hat{\beta}+1.96\sigma$, where $\hat{\beta}$ is the OLS or ML estimate for $\beta$. This nonetheless ignores the parameter uncertainty in estimating $\beta$. This is not a problem if we are just interested in interval consistency, but as noted in Romano et al., (2019), replacing $\beta$ with $\hat{\beta}$ “is not guaranteed to satisfy the coverage statement (…) when $C(X_{n+1})$ is replaced by the estimated interval $\hat{C}(X_{n+1})$.” To address this shortcoming, Romano et al., (2019) employ conformal inference and modify the interval $\hat{C}(X_{n+1})=[\hat{\tau}_{0}(X_{n+1}),\hat{\tau}_{1}(X_{n+1})]$ by adjusting the endpoints to $\hat{\tau}_{0}(X_{n+1})-\vartheta_{1-\alpha}$ and $\hat{\tau}_{1}(X_{n+1})+\vartheta_{1-\alpha}$. Here, $\vartheta_{1-\alpha}$ is the $(1-\alpha)(1+1/n_{2})$-th empirical quantile of the set of conformity scores $\left\{\max\{\hat{\tau}_{0}(X_{i})-Y_{i},Y_{i}+\hat{\tau}_{1}(X_{i})\}:i\in\mathcal{I}_{2}\right\}$ estimated in a calibration sub-sample $\mathcal{I}_{2}$ with sample size $n_{2}$. This corrects for under- or over-coverage and allows for the construction of prediction intervals satisfying desired coverage levels in finite samples using quantile estimates. Our method generalises this idea to interval-valued outcomes. Further details on the conformal inference procedure are provided in Section 4. We also note that although we can construct valid prediction interval with quantile estimates, such intervals are not necessarily minimal in volume, and it is not applicable the oracle prediction set consists of multiple disjoint intervals. Our method addresses these limitations and offers a complementary perspective to traditional quantile-based methods.
@@
When $Y$ is interval-valued and without parametric assumptions, we use recent developments on conformal inference to show formally that a particularly constructed predictive set $\tilde{C}_{n}$ has finite sample coverage property while also enjoying the asymptotic property of being consistent for the oracle prediction set under the partially identified true population distribution. Our conformal procedure designed when $Y$ is interval-valued in turn requires the definition of particular score functions to handle this interval censoring. This score is similar to ones used in the conformal literature but is here modified to handle censoring. In particular, the score function we use is negative when the constructed prediction set contains the interval outcomes, but is otherwise positive, and as a result accounts for both over- and under-coverage as in Romano et al., (2019).
@@
In order to obtain the proposed conformal prediction set, we first characterise the oracle prediction sets under interval censoring and provide a feasible optimisation problem that relates the oracle prediction sets to the conditional distribution of the observed upper and lower brackets, $(Y^{L},Y^{U})$. We use a kernel based conditional distribution function estimator as a basis for the construction of a consistent set estimator for the oracle prediction set. The oracle prediction set may be the union of a collection of disjoint intervals in cases for instance where the prediction density is multimodal. We provide an approach that allows for the set estimator to be a union of disjoint intervals with the prescribed coverage. In cases with a multimodal density, a union of intervals can have a smaller volume than one interval.
@@
We then adapt recent work on conformal inference to propose a conformal prediction set using a sample of exchangeable data. This proposed estimator is shown to be both consistent and have guarantees for finite sample coverage. As a competing alternative prediction set estimator, we show in the appendix that a conformal quantile regression method provides a valid prediction set for interval valued outcomes which can be easier to implement in practice, but is in general not efficient in terms of the volume of the prediction set. In addition, quantile regression based methods are not suitable when the prediction set is not a single interval.
@@
Interval censoring may arise from fixed bracketing and random censoring from “unfolding brackets” or survey schemes, see Heeringa and Suzman, (1995) and Moore and Loomis, (2001). See also the bracketing approach to outcomes using the Health and Retirement Survey in Manski and Tamer, (2002). We thus illustrate our construction using a set of Monte Carlo experiments, where we consider both fixed and random censoring. Our proposed prediction set is shown to have adequate finite sample coverage properties and to have smaller volume compared with the alternative prediction sets constructed with quantile regression methods. In addition, we construct prediction sets using the Current Population Survey (CPS) data. Among people who report their income, a significant fraction (more than 20%) of the observations are interval valued. Interval valued observations are sometimes discarded or handled by imputation methods that rely on assumptions such as missing at random. Our proposed prediction set is robust in the sense that we do not require a specification of the censoring mechanism and we demonstrate using a hold-out subset of the CPS data that our prediction set has better finite sample coverage of the interval valued incomes compared with prediction sets constructed with hot-deck imputation in place. Naturally, without relying on imputation algorithms, our oracle prediction sets are wider than ones obtained with uncensored outcomes, but our constructed sets are more robust and hence more trustworthy.
@@
The paper combines insights from the partial identification literature and conformal inference, which has recently been gaining significant interest. The partial identification literature (see Manski, (2003)) takes the view that one should try and make inferences under minimal assumptions, even if these assumptions do not allow for point identification of the parameter of interest. The motivation for such a view is that inferences are less credible and conclusions are suspect when these rely on non-testable or non credible assumptions. In our context for example, if we are interested in $\beta=E[Y]$ and we observe $[Y^{L},Y^{U}]$ rather than $Y$, then a partial identification approach would just say that, given the data, the identified set for $\beta$ (or all we can learn about $\beta$ given the data) is the interval $\left[E[Y_{0}],E[Y_{1}]\right].$ Imputation based inference uses models to predict a $Y^{P}$ within $[Y^{L},Y^{U}]$ and takes $E[Y^{P}]$ as a proxy for $\beta=E[Y]$. The credibility of this exercise depends on the imputation model, i.e. the set of assumptions it relies on. In the context of interval data, Manski and Tamer, (2002) study identification of linear regression coefficients when the data (on regressors or outcomes) is interval valued using a partial identification approach. See Tamer, (2010), Molinari, (2020), and Kline and Tamer, (2023) for reviews on the developments in the partial identification literature.
@@
In order to obtain a prediction interval that has finite-sample coverage guarantees, we use conformal inference, which is a distribution-free method that builds upon the idea of exchangeability and permutation tests (Vovk et al., (2005), Vovk et al., (2009)). The idea of conformal inference has become increasingly popular in the machine learning literature since it requires no assumptions on the underlying distribution other than the observations being exchangeable and it also allows the use of any prediction protocol that treats the data symmetrically and hence preserves exchangeability. (Barber et al., (2023) discusses the validity of conformal inference when the data is not i.i.d. and the predictor is not symmetric.) Conformal inference has been applied to obtain valid distribution-free prediction sets (Lei et al.,, 2013), in non-parametric regression (Lei and Wasserman,, 2014) and in high-dimensional settings (Lei et al.,, 2018). Romano et al., (2019) considers conformal quantile regression to construct prediction intervals instead of conditional mean regression. Sadinle et al., (2019) proposes a way to construct set-valued classifiers. Chernozhukov et al., (2021) considers construction of valid prediction intervals with distributional regression method. We add to this literature by allowing the outcome observations to be interval-censored or interval-valued. Relatedly,Candès et al., (2023) considers conformal inference for survival analysis. Using our notation, this amounts to observing either $Y_{i}$ or the interval $(Y_{i}^{L},\infty)$ for each individual $i$. They propose a weighted conformal inference procedure (Tibshirani et al.,, 2019) under the assumption that censoring is conditionally independent of her outcome given predictors $X_{i}$. We take a different approach and ask how much information we can extract from a sample with interval-valued outcomes. For example, our procedure allows for the case when all outcomes are interval censored (i.e., $Y_{i}$ is never observed) and the procedure uses the information from the upper bracket $Y^{U}_{i}$ when it is not $\infty$. Also, our procedure does not require the conditional independence assumption which can be violated in many empirical contexts.
@@
There are several related areas connected to our theoretical development. The classical way to construct estimators for the oracle prediction set relies on the level set estimation, (see, for example, Wilks, (1941), Polonik, (1995) and Samworth and Wand, (2010)). The uniform consistency of the nonparametric estimator of the conditional distribution is an extension of classical results on the uniform convergence rate of the kernel-based density and regression estimators such as Einmahl and Mason, (2000), Gine, (2002) and Hansen, (2008). This paper is also related to the literature on quantile regression (see Koenker, (2017) for a review).
@@
This paper is organised as follows. In Section 2, we provide the setup and definitions of optimal sets and approaches to obtaining estimators of such sets with censoring. Section 3 provides a study of consistency of such sets using kernel estimators of the conditional distribution. Section 4 provides results on finite sample coverage using a conformal inference procedure under interval censoring. Section 5 provides some Monte Carlo evidence and Section 6 applies our methods to US CPS Data. Section 7 concludes. The proofs for the theorems in the main text as well as some further results and comments are collected in the Appendix.
@@ 2 Setup and definitions
Suppose there is a random sample $\bar{\mathcal{D}}=\{\bar{Z}_{i}=(X_{i},Y_{i},Y_{i}^{L},Y_{i}^{U})\in\mathbb{R}^{d+3}:i\in\mathcal{I}\}$ with sample size $\left\lvert\mathcal{I}\right\rvert=n$ drawn independently from a joint distribution $P$. We only observe the sub-vector $Z_{i}=(X_{i},Y_{i}^{L},Y_{i}^{U})$, and $Y_{i}$ is not directly observed but assumed to satisfy the condition $Y_{i}^{L}\leq Y_{i}\leq Y_{i}^{U}$. For each $i\in\mathcal{I}$, the predictor $X_{i}$ is a $d$-dimensional random vector with support $\mathcal{X}\subset\mathbb{R}^{d}$. The latent outcome $Y_{i}$ and the observed lower and upper bounds $Y_{i}^{L}$ and $Y_{i}^{U}$ are one-dimensional. Let $\bar{Z}=(X,Y,Y^{L},Y^{U})$ denote a generic sample from the joint distribution $P$. We will use notation such as $P_{Y}$ to denote the marginal distribution of $Y$, and $P_{Y\mid X}$ to denote the distribution of $Y$ conditional on $X$. Note here that the joint distribution $P_{X,Y^{L},Y^{U}}$ of $(X,Y^{L},Y^{U})$ is identified given the observed sample $\mathcal{D}=\left\{Z_{i}=(X_{i},Y_{i}^{L},Y_{i}^{U})\in\mathbb{R}^{d+2}:i\in\mathcal{I}\right\}$, but the joint distribution $P_{X,Y}$ of predictors $X$ and latent outcome $Y$ is not point identified given the observed sample $\mathcal{D}$ without restrictions on the censoring mechanism.
A prediction set is defined as a set $C\subset\mathcal{X}\times\mathcal{Y}$, with the interpretation that, given a realisation of the predictor $X=x$, we predict that $Y$ is likely to fall inside the section $C(x)=\left\{y:(x,y)\in C\right\}$. In the absence of interval censoring, estimation of a prediction set that satisfies certain optimality criteria (i.e. the minimal prediction set with coverage guarantees) has been studied in the literature; see Lei and Wasserman, (2014), Barber et al., (2021) among others. As we discuss later in this section, the oracle prediction set without interval censoring is related to the upper level set of the density function of $Y$ given $X$ under point identification.
@@
There are two criteria for an oracle prediction set: validity and efficiency (see Vovk et al.,, 2009; Lei et al.,, 2013). Let $\mathcal{P}_{I}$ be the identified set of the joint distribution of $(X,Y)$ under censoring. That is, $\mathcal{P}_{I}$ is the set of joint distributions of $(X,Y)$ that are compatible with the joint distribution $P_{X,Y^{L},Y^{U}}$ and some censoring mechanism. A prediction set $C$ is valid if it satisfies the following coverage properties for a new sample $(X_{n+1},Y_{n+1})$ drawn from the same distribution $P$ at a given miscoverage level $\alpha$.
@@
When the model is point identified, $\mathcal{P}_{I}=\left\{P_{X,Y}\right\}$ is a singleton set, the conditions in Equations (2) and (3) reduce to the standard conditions of validity, see for example, Lei and Wasserman, (2014). We will simply refer to these two conditions as marginal validity and conditional validity henceforth.
It is natural to require a prediction set to be valid, so that we have confidence it will contain the true value of the outcome with a certain probability. However, the requirement of validity by itself is vacuous. For example, the prediction set $C$ such that $C(x)=(-\infty,\infty)$, for all $x\in\mathcal{X}$ is technically valid at any miscoverage level but it is also uninformative. Therefore, our target parameter of interest is an oracle prediction set, denoted by $C^{*}_{\mathcal{P}_{I}}$ which is defined as the minimal-volume prediction set that satisfies the validity conditions in Definition 2.1. For simplicity, we will omit the subscript and write $C^{*}=C^{*}_{\mathcal{P}_{I}}$ whenever the dependence on the identified set ${\mathcal{P}_{I}}$ is clear. Since the minimal-volume prediction set is intuitively the most informative among all valid prediction sets, this property is also called efficiency. [^1] This notion of efficiency, (see Vovk et al.,, 2009; Lei et al.,, 2013; Lei and Wasserman,, 2014) differs from the usual definition, where an estimator is considered efficient if it has the smallest variance within a class of estimators. However, the concepts are related, as an estimator with smaller variance generally leads to the construction of tighter and more informative confidence intervals.
For interpretability, we will focus on the prediction sets such that $C(x)$ takes the form of a union of a (potentially unknown) number of disjoint intervals for each $x\in\mathcal{X}$. Specifically, we assume that $C(x)\in\mathcal{C}$, $P_{X}$-almost surely, where
@@
where $\mu$ denotes the Lebesgue measure. Since conditional validity implies marginal validity, the oracle prediction set will also be marginally valid under partial identification.
Before proposing a feasible estimation procedure for the oracle prediction set $C^{*}$, we explain why the existing estimation methods are not directly applicable for prediction sets in the case of interval outcome observations. This analysis highlights the challenges posed by partial identification and the difference between our approach and the classical level set method studied in the literature.